OpenStudy (anonymous):

Solve for x and y 12sinx+5cosx = 2y^2-8y+21

ganeshie8 (ganeshie8):

you can rewrite left hand side as \[\sf 12\sin x + 5 \cos x = \langle 5, 12\rangle \cdot \langle \cos x, \sin x\rangle = \sqrt{5^2+12^2}\cos\left(x-\arctan\left( \frac{12}{5}\right)\right)\]

ganeshie8 (ganeshie8):

try completing the square for right hand side

ganeshie8 (ganeshie8):

\[\sf 2y^2 - 8y+21 = 2(y-2)^2 + 13\]

ganeshie8 (ganeshie8):

\[\sf 13\cos\left(x-\arctan\left( \frac{12}{5}\right)\right) = 2(y-2)^2+13 \] Notice that the maximum value of left hand side is 13 and the minimum value of right hand side is 13. What can you conclude from that ?

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

yw! so whats your solution for x and y ?